[figure: 9.7 cm]

Transport properties of lanthanum substituted CeF e_{4}Sb_{12}

R. Viennois ^{a}, F. Terki ^{a}, A. Errebbahi ^{a}, S. Charar ^{a}, D. Ravot ^{b}, C.

Tedenac ^{b}, P. Haen ^{c}, C. Sekine ^{d}

a Groupe d’Etude des Semiconducteurs, UMR 5650 Université Montpellier II, Place

Eugène Bataillon 34095 cedex 5 Montpellier, France.

b Laboratoire de Physico-Chimie de la matière Condensée, UMR 5617, Université

Montpellier II, Place Eugène Bataillon 34095 cedex 5 Montpellier, France. c Centre de Recherche sur les Très Basses Températures, 25 Avenue des Martyrs,

BP 166, 38042 Grenoble Cedex 09 France. d Departement of Electrical and Electronic Engineering Muroran Institute of Technology 27-1. Mizumoto-cho, Muroran, Hokaido 050-8585, Japan

The skutterudite CeF e_{4}Sb_{12} exhibits a quite large low temperature Sommerfeld coefficient. However, no consistent explanation of the temperature dependences of its magnetic and transport properties has been given so far. To add to the confusion, previous results show spurious anomalies which have been related to foreign phases. We present here new measurements performed on good polycrystals: resistivity of Ce_{1−z}La_{z}F e_{4}Sb_{12} alloys and magneto-transport study of CeF e_{4}Sb_{12}. The resistivity of CeF e_{4}Sb_{12} shows a huge decrease at T < 150K which previous authors have attributed to the onset of the coherent electron scattering from rare-earth sublattice. However, according to magnetic susceptibility results, the crystal field effects could contribute to the splitting of the Ce^{3+} ground state in the resistivity behavior, as usual for a Kondo lattice compound. This idea will be discussed in view of our magnetic and magneto-transport experiments

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : poster

Corresponding author : F. Terki

Address for correspondence : Groupe d’Etude des Semiconducteurs, UMR 5650 Université Montpellier II, CC074, Place Eugène Bataillon 34095 cedex 5 Montpellier, France

Email address : ferial@ges.univ-montp2.fr